what is the length of the line segment that joins the points whose coordinates are (4,7) and (-3,5)

Answer by stanbon(75887)

what is the length of the line segment that joins the points whose coordinates are (4,7) and (-3,5)
-----
d = sqrt[(7-5)^2+(4--3)^2] = sqrt[4+49] = sqrt(53)
================
Cheers,
Stan H.
=============